Grade 9General Physics

A particle travels in a circular path of radius 0.2 meter with a constant energy of 4J what is the net force on this particle????

Asma
7 Years agoGrade 9

1 Answer

Arun
7 Years ago
KE = 1/2 mv^2 and we know from studying centripetal forces that F=mv^2/r 

so mv^2 = 2 KE, therefore, mv^2=8J 

F = mv^2/r = 8J/0.2m = 40N